BOOK REVIEW

Thomas Dence

Sally McClain, Navajo Weapon: The Navajo Code Talkers, Rio Nuevo Publishers, Tucson, AZ, 2001, 304 pp. $16.95.

During the summer of 2002 the Nicholas Cage movie “The Windtalkers” appeared on the cinema screen. This nonfiction movie chronicled an episode in United States history that occurred in the Pacific during World War II.

It was here that the Marine Corps made initial and significant use of the Navajo Indians to handle radio transmissions during manuevers and, especially, battles. This was not the first time the U. S. Military made similar use of native Americans during wartime. During World War I, for example, a wide variety of American Indians were used to transmit messages via field phones or radios in their native tongue to confuse on–listening enemies. The Choctaw soldiers used a code created from their language to baffle the German cryptologists during Germany’s final attacks in France. Other tribes represented by the Cheyenne, Comanche, Cherokee, Osage, and Sioux also parlayed their languages to aide the Allied cause.

Whereas “The Windtalkers” was extremely weak of describing cryptographic details, which probably contributed to the movie’s downfall, this novel by Sally McClain goes into much detail. We learn how the idea for the use of the Navajos was germinated, and how an initial batch of 29 Indians, mainly from the Southwest, were recruited and put through boot camp at San Diego. They then moved to Camp Elliott where they worked in earnest on the “code” for the military, and the Navajos had to work together to correlate military terms with Navajo vocabulary. For instance, the term “Commanding General” gets translated to “War Chief” and is pronounced “Bih–keh–he”, while “dive bomber” is translated to “chicken hawk”, and “tank” is translated as “tortoise.”

The Navajo code talkers got their first taste of battle at Guadalcanal, with subsequent encounters at Bougainville, New Britain, Saipan, Tinian, Guam, Peleliu, and Iwo Jima. It was on Iwo Jima where the Navajo helped celebrate the securing of Mt. Suribachi in the famous flag raising photograph of Joe Rosenthal. We learn how proud the Navajos were to serve their country, and how they continued to look upon their wartime activities as time well–spent. Immediately after the war, and in fact for about 20 years, the government wanted all knowledge of their involvement to be hush–hush just in case they needed to be called back to active duty.

It is clear that Sally McClain has done a terrific job with her homework. She searched through a multitude of documents and conducted many personal interviews to gather important details of this story, along with quite a few occurrences of a personal touch involving the Navajo code talkers. Her writing style is clean, orderly, thorough, and easy to follow.

As an interesting sidenote, three of the original 29 Navajo code talkers had the same last name of Begay, which happens to be a common Navajo name. Many more Begays were to follow during the next 200 recruits, with one in particular being Notah Begay. His great grandson, of the same name, would make national headlines in 2000 as being one of the top players on the PGA golf tour, winning several tournaments during the year. Furthermore, this grandson had sharpened his golfing skills in college at Stanford University where he teamed with a young fellow from Southern California named Tiger Woods.

Thomas Dence Department of Mathematics and Computer Science Ashland University Ashland, OH 44805